OpenStudy (anonymous):

write an equation in slope-intercept form of the line that passes through the point (-4,1) and has a slope of -3/2 PLEAASSEE HELLLPP !!!???

OpenStudy (anonymous):

y-y1=m(x-x1) \[y-1=-\frac{ 3 }{ 2 }\left( x-\left( -4 \right) \right)\] 2y-2=-3(x+4) 2y-2=-3x-12 2y=-3x-12+2 2y=-3x-10 \[y=-\frac{ 3 }{ 2 }x-\frac{ 10 }{ 2 }=-\frac{ 3 }{ 2 }x-5\]

OpenStudy (anonymous):

Write the equation of the line in point-slope form for the line through the given point (4,-6) and with m=3/5 do you have the answer to that one please?